# Adversarial Flow Models

### v1

@@ 3.3 Multi-step Adversarial Flow Models
Adversarial flow models can be generalized to multi-step generation and allow the model to jump between arbitrary timesteps on the probability flow. We introduce an interpolation function of the same form as flow-matching models:
@@
where $t\in[0,1]$. Although $A(t)$ and $B(t)$ can be any continuous monotonic functions satisfying $x_{0}=x,x_{1}=z$, we adopt linear interpolation, where A(t)=1tA(t)=1-t and $B(t)=t$, to match the $\mathrm{W}^{2}_{2}$ transport behavior.
We change the generator to accept additional source timestep $s$ and target timestep $t$ as $G(x_{s},s,t)$, and the discriminator to accept additional target timestep as $D(x_{t},t)$. During training, $x_{s},x_{t}$ are obtained by interpolating randomly sampled $x,z$. The adversarial loss is extended as:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{x_{s},x_{t},s,t}\left[f(D(x_{t},t)-D(G(x_{s},s,t),t))\right],$ (16)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{x_{s},x_{t},s,t}\left[f(D(G(x_{s},s,t),t)-D(x_{t},t))\right].$ (17)
The $R_{1}$ and $R_{2}$ gradient penalties are substituted accordingly. We omit the approximation form for simplicity:
$\mathcal{L}^{D}_{\mathrm{r1}}$ $=\mathbb{E}_{x_{t},s,t}\left[w(s,t)\ \|\triangledown_{x_{t}}D(x_{t},t)\|^{2}_{2}\right],$ (18)
$\mathcal{L}^{D}_{\mathrm{r2}}$ $=\mathbb{E}_{x_{s},s,t}\left[w(s,t)\ \|\triangledown_{G(x_{s},s,t)}D(G(x_{s},s,t),t)\|^{2}_{2}\right].$ (19)
In flow-matching models with linear interpolation, the transport plan between any source and target timestep’s marginal distributions also minimizes the $\mathrm{W}_{2}^{2}$ distance, so the optimal transport loss is generalized as:
$\small\mathcal{L}^{G}_{\mathrm{ot}}=\mathbb{E}_{x_{s},s,t}\left[\frac{1}{n}\cdot\frac{1}{w(s,t)}\cdot\|G(x_{s},s,t)-x_{s}\|^{2}_{2}\right].$ (20)
@@
where $\delta=0.001$ is set for numerical stability.
During training, the timesteps can be sampled as continuous values: $s\sim\mathcal{U}(0,1),t\sim\mathcal{U}(0,s)$. In this case, the model is trained to support transporting between any arbitrary timesteps $t<s$ on the probability flow for generation. When $s$ and $t$ are close, the model behaves like flow-matching models. When they are far, the model behaves like trajectory models. Since $G$ directly learns the target distribution through $D$ without needing consistency propagation, the model can alternatively be trained to only learn a discrete set of timesteps that is needed for a specific few-step inference setting, where single-step generation is just a special case. This saves model capacity and training iterations. Our framework unifies single-step and flow generation under the same adversarial training setting.
@@ 3.4 Discriminator Formulation
It is important not to condition the discriminator on the source sample. Specifically, formulating the discriminator as $D(x,\underline{z})$ for single-step, or $D(x_{t},t,\underline{x_{s},s})$ for multi-step, is incorrect. This is because at training, $x$ and $z$ are randomly sampled and paired. This formulation incorrectly tells $D$ that $z$ should be paired with all $x$, but $G$ can only produce a single mapping, and it is impossible to achieve such a goal, so the training will oscillate or diverge.
When searching for hyperparameters, a complication we have encountered is gradient magnitude. Specifically, the $G$ objective consists of both the adversarial and the optimal transport losses. However, the adversarial loss received from $D$ may have varying magnitudes, influenced by the $D$ architecture, its weight initialization, the gradient penalty strength, and the adversarial objective. This creates difficulties in finding $\lambda_{\mathrm{ot}}$ across model sizes.
Formally, we break down the gradient of the generator loss with respect to the generator network parameter $\theta$ following the chain rule:
@@
The boxed terms are the gradient passed down from $D$ that has varying magnitudes. In traditional GANs, only the adversarial loss is used, and adaptive optimizers rescale the magnitude per parameter so $G$ is invariant to the absolute scale of the gradient. However, in our case, the magnitude matters because it determines the ratio between the adversarial and the optimal transport losses.
Therefore, we propose a gradient normalization technique. Specifically, we change the formulation to $D(\phi(G(z)))$, where $\phi$ is an identity operator in the forward pass but rescales the gradient in the backward pass:
@@
The operator $\phi$ tracks the exponential moving average (EMA) of the gradient norm, normalizes the gradient by the average norm, and rescales by $\frac{1}{\sqrt{n}}$ where $n$ is the data dimension. It can be seen as an extension of Adam [36] into the backward path, so we use the same $\beta_{2}$ as in Adam for the EMA decay. After normalizing the adversarial gradient to a unit scale, we can find $\lambda_{\mathrm{ot}}$ that works well across model sizes.
3.5 Better Distribution Matching
Adversarial flow models enable the complete disentanglement of distribution-matching and optimal-transport objectives. Although the optimal-transport objective matches flow matching, the distribution-matching objective does not.
Flow matching’s squared distance loss objective minimizes the Euclidean distance in the $\mathbb{R}^{n}$ space instead of the semantic distance on the data manifold [43]. As a result, in high-dimensional space, new data samples generated by the model through generalization appear element-wise blended instead of semantically morphed from the training data. This is why flow-matching models without guidance generate perceptually out-of-distribution samples. The effect of classifier-free guidance (CFG) [19] is not only low-temperature sampling, but also perceptual guidance [43]. However, CFG entangles perceptual quality with strong conditional alignment. This limits flow-matching models from representing the full data distribution and can make samples unnatural and canonical [45, 33]. Consistency-based methods share the same pathology. Latest works have explored training on more representational latent spaces [103, 41] and using other guidance types [33, 34, 93] to counter this congenital defect. (More in Appendix E.)
Adversarial training is different because of the introduction of a learnable $D$. The distance between real and generated distributions is measured through the semantic network, which captures the perceptual and semantic distance [102] instead of the Euclidean distance. Our experiments find that our model can outcompete flow matching in the guidance-free setting.
@@ 3.6 Connections to Guidance
Having control over sampling temperature and conditional alignment is still a desirable property. We show that guidance can be incorporated into adversarial flow models. We use classifier guidance (CG) [9] for conditional generation as an illustration because of its popularity.
@@
Figure 3: Conditional models trained on 1D Gaussian mixtures. (a) The conditional flow. (b) The effect of CFG on flow-matching models. (c) Adversarial flow models with simple classifier guidance does not work. (d) Adversarial flow models with flow-based classifier guidance match the behavior of CFG.
We visualize an example conditional flow in Fig. 3(a) and the effect of CFG on flow-matching models in Fig. 3(b). Prior adversarial works [73, 26] introduce a classifier $C(x,c)$ that predicts $p(c|x)$, and train $G$ with an additional loss to maximize the classification probability:
$\small\mathcal{L}^{G}_{\textrm{cg}}=\mathbb{E}_{z,c}\left[-C(G(z,c),c)\right],$ (24)
$\small\mathcal{L}^{G}_{\mathrm{cAF}}=\mathcal{L}^{G}_{\mathrm{adv}}+\lambda_{\mathrm{ot}}\mathcal{L}^{G}_{\mathrm{ot}}+\boxed{\lambda_{\mathrm{cg}}\mathcal{L}^{G}_{\mathrm{cg}}}.$ (25)
However, as Fig. 3(c) shows, the guided model yields almost identical results to the original. This is because in this particular example, the classes are well separated, so the classifier has a clear decision boundary and yields no gradient. This example illustrates the important difference compared to CFG, whose transport is influenced by guidance gradients accumulated along the flow instead of at a single timestep. The classifier gradients exist on higher timesteps due to the diffusion of class boundaries by interpolation.
To obtain the accumulated guidance gradient along the flow, even for single-step training and generation, we switch to a time-conditioned classifier $C(x_{t},t^{\prime},c)$ that predicts $p(c|x_{t^{\prime}})$ on a probability flow. During training, $G$,$D$ are still trained for a single step only. The generated samples $G(z,c)$ are interpolated to random timesteps $t^{\prime}\sim\mathcal{U}(0,1)$ before feeding to the classifier:
@@
This maximizes the expectation of classification on all timesteps, giving us similar results to CFG as shown in Fig. 3(d). The timestep $t^{\prime}$ can be sampled from a custom range, which corresponds to performing CFG on selected timesteps. Eq. 24 is a special case of Eq. 26 where $t^{\prime}$ is always 0. The hyperparameters, i.e. the scale $\lambda_{\textrm{cg}}$ and the range of $t^{\prime}$, can be optionally amortized into $G$ to allow inference-time adjustment.
For clarity, our experiment trains $C$ offline as an independent network. If adversarial flow is used for post-training an existing flow-matching model $v$, the gradient of an implicit classifier can be derived from $v$ following CFG:
$\small\triangledown_{x_{t}}C(x_{t},t,c)=v(x_{t},t)-v(x_{t},t,c),$ (27)
$\small\mathcal{L}^{G}_{\textrm{cfg}}=\mathbb{E}_{z,c,z^{\prime},t^{\prime}}\left[-\frac{1}{n}\sum G(z,c)\,\triangledown_{(\cdot)}C(\cdot,t^{\prime},c)\right],$ (28)
where $(\cdot)$ is short for $\textit{interp}(G(z,c),z^{\prime},t^{\prime})$.
@@ 3.7 Model Architecture
We parameterize the single-step or multi-step generator $G$ using a neural network $g$. We find that it can work equally well by either formulating it directly:
@@
The latter is closely related to the velocity prediction formulation of existing flow-matching models and consistency-based models. To demonstrate the feasibility of both, we train our single-step models using the direct formulation and multi-step models using the residual formulation. We parameterize $D$ directly using a neural network $d$.
Both the $g$ and $d$ networks use standard diffusion transformer architecture (DiT) [62]. For single-step models, the timestep projection is removed. For multi-step models with fixed discretizations, one timestep projection is used. For any-step models, two timestep projections are used in $g$. The condition is injected through modulation for both $g$ and $d$ following the original DiT. Our $d$ is nearly identical to $g$, except for the addition of a learnable [CLS] token prepended to the input. The [CLS] token is used to produce the logit through a final LayerNorm [2] and a linear projection. Our architecture has minimal modifications to the original DiT.
@@ 3.8 Deep Model Architecture
Figure 4: Deep architecture through transformer block repetition. Extra-deep models are trained end-to-end using single-step objective without any intermediate supervision.
Prior research [47] indicates that the effective model depth is critical for capturing the non-linear transformations required for generating high-quality samples, and the lack thereof is the primary reason causing artifacts on single-step models. We experiment with end-to-end training of extra-deep single-step models. In theory, extra-deep single-step models can outperform their multi-step counterparts because the models can pass hidden states end-to-end without projecting to and reinterpreting from the data space, there is no manual definition of timestep discretizations, and the network is trained without teacher-forcing.
As illustrated in Fig. 4, our extra-deep models use transformer block repetition [8]. The hidden state after the initial pass is recycled. Timestep-like embedding is still given to the transformer blocks only to differentiate repetition iterations, but the whole network is trained end-to-end for single-forward generation without any intermediate supervision. This design allows us to exactly match the number of parameters and model behavior with the multi-step counterpart for comparison.
3.9 Techniques for Adversarial Training
We share additional training techniques that are, though imperfect, very effective. They address the remaining challenges of adversarial training, not introduced by our adversarial flow models.
On optimization, minimax optimization with gradient descent is prone to oscillation. The equilibrium is better achieved through the weight average than at the last iterate [7]. We find optimistic optimizer [7] and asymmetrical learning rates [18] ineffective. Our approach is to keep an EMA of $G$ [95], which consistently outperforms the online model by a large margin. More importantly, once the performance of the EMA model plateaus toward the end of the training, we replace online $G$ with the EMA weights and repeat automatically after every epoch onward. Learning rate is also decreased during this phase. We find this simple technique very effective for approaching peak performance.
On training dynamics, vanilla adversarial training needs $G$ and $D$ to learn at a synchronized pace. If $G$ learns too slowly, $D$ will separate the real and the generated distributions perfectly, and the gradient vanishes. To address the vanishing gradient problem, WGAN [1] is proposed but requires a K-Lipschitz $D$. DiffusionGAN [90] projects $D$ onto a flow in the same spirit as our approach with guidance, but this only guarantees support up to the signal-to-noise ratio that $G$ can perfectly fool $D$ under each’s capacity. Discriminator augmentation (DA) [30] is another approach to increase the distribution overlap used by recent GANs [21, 23]. However, we find that the choice of augmentation may implicitly inject inductive biases. For example, prior work finds that affine transforms outperform other distortions on image data [30] because it implicitly encourages $D$ to recognize affine transforms as a more acceptable generalization for image generation. The last approach is to simply reload $D$ from an earlier checkpoint to reset the pace. Our experiment finds that reloading $D$ performs surprisingly well while DiffusionGAN has limited effectiveness. Therefore, we reload $D$ when training stalls for the no-guidance setting to avoid introducing any additional biases, and additionally use DA for the guidance setting.
@@ 4.1 Setups and Training Details
We train on class-conditional ImageNet [70] generation to compare with prior works. We follow the protocols to resize the images to $\times256$ resolution and augment with horizontal flips. We use pre-trained variational autoencoder (VAE) [^1] https://huggingface.co/stabilityai/sd-vae-ft-mse [68] and train the models in $32\times 32\times 4$ latent space. Evaluations use Fréchet Inception Distance on 50k class-balanced samples (FID-50k) [18] against the entire train set.
@@
We use a learning rate of $1\times 10^{-4}$ and a batch size of 256 consistent with prior works [62, 12, 14]. We use AdamW [51] optimizer with $\beta_{1}=0,\beta_{2}=0.9$ [47]. We set weight decay to 0.01 and EMA decay to 0.9999. We follow MeanFlow’s definition of model sizes: B/2, M/2, L/2, XL/2, where 2 denotes the patch size. $G$ and $D$ always use models of the same size. We use separate dataloaders for $G$ and $D$. Epochs are measured as the number of images seen by $G$. Since different models reach peak FID at different epochs, we report the earliest epoch that reaches the best FID. Additional training details are in the appendix.
@@
We train the model without guidance till the best FID, and then continue training it with guidance. Our classifier is trained from scratch on ImageNet using the cross-entropy objective for 30 epochs. It uses the same B/2 transformer in the latent space. We do not amortize the scale of CG into models for a fair comparison with prior works.
@@
We only repeat depth on $G$ while keeping $D$ normal depth. The learning rate of $G$ is reduced by the factor of repetition. Extra-deep models are trained end-to-end from scratch following the single-step objective.

### v2

@@ 3.3 Multi-step Adversarial Flow Models
Adversarial flow models can be generalized to multi-step generation, allowing the model to transport between arbitrary timesteps along the probability flow. We introduce an interpolation function of the same form used in flow-matching models:
@@
where $t\in[0,1]$. For simplicity, we adopt linear interpolation, with A(t)=1tA(t)=1-t and $B(t)=t$.
We modify the generator to accept an additional source timestep $s$ and target timestep $t$, yielding $G(x_{s},s,t)$, and we modify the discriminator to accept the target timestep, yielding $D(x_{t},t)$. During training, $x_{s}$ and $x_{t}$ are obtained by interpolating independently sampled $x$ and $z$. The adversarial loss then becomes:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{x,z,s,t}\left[f(D(x_{t},t),D(G(x_{s},s,t),t))\right],$ (16)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{x,z,s,t}\left[f(D(G(x_{s},s,t),t),D(x_{t},t))\right].$ (17)
The $R_{1}$ and $R_{2}$ gradient penalties are modified accordingly. We omit their approximate forms for simplicity:
$\mathcal{L}^{D}_{\mathrm{r1}}$ $=\mathbb{E}_{x,z,s,t}\left[w(s,t)\ \|\nabla_{x_{t}}D(x_{t},t)\|^{2}_{2}\right],$ (18)
$\mathcal{L}^{D}_{\mathrm{r2}}$ $=\mathbb{E}_{x,z,s,t}\left[w(s,t)\ \|\nabla_{G(x_{s},s,t)}D(G(x_{s},s,t),t)\|^{2}_{2}\right].$ (19)
In the multi-step setting, Brenier’s theorem still applies because the source distribution remains absolutely continuous, as the interpolation process is equivalent to convolution with a Gaussian. The quadratic optimal transport loss can be generalized as follows:
$\small\mathcal{L}^{G}_{\mathrm{ot}}=\mathbb{E}_{x,z,s,t}\left[\frac{1}{n}\cdot\frac{1}{w(s,t)}\cdot\|G(x_{s},s,t)-x_{s}\|^{2}_{2}\right].$ (20)
@@
where $\delta=0.001$ is chosen for numerical stability.
During training, the timesteps can be sampled as $s\sim\mathcal{U}(0,1)$ and $t\sim\mathcal{U}(0,s)$. In this case, the model is trained to support transport between arbitrary timesteps with $t<s$ along the probability flow at generation time. When $s$ and $t$ are close, the model behaves like a flow-matching model. When they are far apart, the model behaves like a trajectory model. Since $G$ directly learns the target distribution through $D$ without requiring consistency propagation, the model can alternatively be trained only on the discrete set of timesteps needed for a specific few-step inference setting, of which single-step generation is a special case. This saves model capacity and training iterations. Our framework therefore extends adversarial training from single-step generation to discrete-time flow modeling.
@@ 3.4 Discriminator Formulation
It is important not to condition $D$ on the source sample. Specifically, formulating the discriminator as $D(x,\underline{z})$ for the single-step setting, or as $D(x_{t},t,\underline{x_{s},s})$ for the multi-step setting, is incorrect. This is because, during training, $x$ and $z$ are sampled independently. This formulation incorrectly tells $D$ that $z$ should be paired with every $x$, but $G$ can produce only a single mapping. Since this objective is impossible to satisfy, training will oscillate or diverge.
When searching for hyperparameters, one complication we encounter is gradient magnitude. Specifically, the objective for $G$ consists of both the adversarial loss and the optimal transport loss. However, the adversarial gradient received from $D$ can have varying magnitudes, influenced by the architecture, the weight initialization, and the gradient-penalty strength. This makes it difficult to find a value of $\lambda_{\mathrm{ot}}$ that works across model sizes.
Formally, we decompose the gradient of the generator loss with respect to the generator parameter $\theta$ using the chain rule:
@@
The boxed term is the gradient passed down from $D$, and its magnitude can vary substantially. In traditional GANs, only the adversarial loss is used, and adaptive optimizers rescale the magnitude of each parameter update, making $G$ largely invariant to the absolute gradient scale. In our case, however, the magnitude matters because it determines the ratio between the adversarial and optimal transport losses.
We therefore propose a gradient-normalization technique. Specifically, we change the formulation to $D(\phi(G(z)))$, where $\phi$ is the identity operator in the forward pass but rescales the gradient in the backward pass:
@@
The operator $\phi$ tracks the exponential moving average (EMA) of the gradient norm, normalizes the gradient by this average norm, and rescales it by $\frac{1}{\sqrt{n}}$, where $n$ is the data dimensionality. It can be viewed as an extension of Adam (Kingma, 2014) to the backward path, so we use the same $\beta_{2}$ as in Adam for the EMA decay. After normalizing the adversarial gradient to a unit scale, we can find a value of $\lambda_{\mathrm{ot}}$ that works well across model sizes.
@@ 3.5 Connections to Guidance
Control over sampling temperature and conditional alignment is a desirable property. We show that guidance can be incorporated into adversarial flow models. As illustrative examples, we consider classifier guidance (CG) (Dhariwal and Nichol, 2021) and classifier-free guidance (CFG) (Ho and Salimans, 2022), given their popularity in conditional generation.
We visualize an example conditional flow in Figure 3(a) and the effect of CFG on flow-matching models in Figure 3(b). Prior adversarial works (Sauer et al., 2022; Kang et al., 2023) introduce a classifier $C(x,c)$ that predicts $p(c|x)$ and train $G$ with an additional loss that maximizes the classification probability:
$\mathcal{L}^{G}_{\textrm{cg}}=\mathbb{E}_{z,c}\left[-C(G(z,c),c)\right],$ (24)
$\mathcal{L}^{G}_{\mathrm{AF}}=\mathcal{L}^{G}_{\mathrm{adv}}+\lambda_{\mathrm{ot}}\mathcal{L}^{G}_{\mathrm{ot}}+\boxed{\lambda_{\mathrm{cg}}\mathcal{L}^{G}_{\mathrm{cg}}}.$ (25)
However, as Figure 3(c) shows, the guided model yields results that are almost identical to those of the original model. This is because, in this particular example, the classes are well separated, so the classifier has a clear decision boundary and yields no gradient. This example illustrates an important difference from CFG, whose transport is influenced by guidance gradients accumulated along the flow rather than at a single timestep. The classifier gradients exist at higher timesteps because interpolation diffuses the class boundaries.
@@
Figure 3: Conditional models trained on a 1D Gaussian mixture. (a) The conditional flow. (b) The effect of CFG on flow-matching models. (c) Adversarial flow models with simple classifier guidance does not work. (d) Adversarial flow models with flow-based classifier guidance match the behavior of CFG.
To obtain the accumulated guidance gradient along the flow, even for single-step training and generation, we switch to a time-conditioned classifier $C(x_{t^{\prime}},t^{\prime},c)$ that predicts $p(c|x_{t^{\prime}})$ on the probability flow. During training, $G$ and $D$ are still trained for a single step only. The generated samples $G(z,c)$ are interpolated to random timesteps $t^{\prime}\sim\mathcal{U}(0,1)$ with independent noise samples $z^{\prime}\sim\mathcal{Z}$ before being fed to the time-conditioned classifier:
@@
This maximizes the expected classification score over all timesteps, giving results similar to CFG, as shown in Figure 3(d). The timestep $t^{\prime}$ can be sampled from a custom range, which corresponds to performing CFG on selected timesteps. Equation 24 is a special case of Equation 26 in which $t^{\prime}$ is always 0. The hyperparameters, i.e., the scale $\lambda_{\textrm{cg}}$ and the range of $t^{\prime}$, can optionally be amortized into $G$ to allow inference-time adjustment.
For clarity, in our experiment, $C$ is trained offline as an independent network. If adversarial flow is used to post-train an existing flow-matching model $v$, the gradient of an implicit classifier can be derived from $v$ following CFG:
$\small\nabla_{x_{t}}C(x_{t},t,c)=v(x_{t},t)-v(x_{t},t,c),$ (27)
$\small\boxed{\mathcal{L}^{G}_{\textrm{cfg}}=\mathbb{E}_{z,c,z^{\prime},t^{\prime}}\left[-\frac{1}{n}G(z,c)^{\top}\nabla_{(\cdot)}C(\cdot,t^{\prime},c)\right],}$ (28)
where $(\cdot)$ is shorthand for $\textit{interp}(G(z,c),z^{\prime},t^{\prime})$. Details and derivations are provided in Appendix F.
3.6 Different Model Generalization
In theory, under infinite capacity and perfect optimization, flow matching converges to the ground-truth probability flow, and adversarial training reaches equilibrium. Both types of models transport samples to the empirical distribution of the training data and therefore overfit by reproducing only the training samples. In practice, however, finite-capacity models learn a generalized distribution. Adversarial training can induce a different generalized distribution.
Specifically, flow matching’s squared $L_{2}$ criterion measures isotropic Euclidean distance rather than semantic distance on the data manifold. Furthermore, flow matching, connected to DDPM (Ho et al., 2020), minimizes forward KL divergence to maximize mode coverage. These properties lead to frequent generation of out-of-distribution samples in the guidance-free setting. In contrast, adversarial training uses a discriminator network as a learned criterion. Prior work suggests that deep networks can better capture manifold structure and serve as a better perceptual metric than Euclidean distance (Zhang et al., 2018). Our choice of adversarial objective is also closer to JS divergence (Goodfellow et al., 2014) and is less sensitive to outliers. We hypothesize that these factors help explain why our models even outperform flow matching in the guidance-free setting.
@@ 3.7 Model Architecture
We parameterize the single-step or multi-step generator $G$ with a neural network $g$. We find that it works equally well when formulated either directly:
@@
The latter is closely related to the velocity-prediction formulation used in existing flow-matching and consistency-based models. To demonstrate the feasibility of both, we train our single-step models using the direct formulation and our multi-step models using the residual formulation. We parameterize $D$ directly with a neural network $d$.
Both $g$ and $d$ use the standard diffusion transformer (DiT) architecture (Peebles and Xie, 2023). For single-step models, the timestep projection is removed. For multi-step models with fixed discretizations, a single timestep projection is used. For any-step models, two timestep projections are used in $g$. The condition is injected through modulation in both $g$ and $d$, following the original DiT. Our discriminator $d$ is nearly identical to $g$, except for the addition of a learnable [CLS] token prepended to the input. The [CLS] token is used to produce the logit through a final LayerNorm (Ba et al., 2016) and a linear projection. Overall, our architecture requires only minimal modifications to the original DiT.
@@ 3.8 Deep Model Architecture
Figure 4: Deep architecture using transformer block repetition. Extra-deep models are trained end-to-end using single-step objective without any intermediate supervision.
Prior research (Lin et al., 2025a) indicates that effective model depth is critical for capturing the nonlinear transformations required to generate high-quality samples, and that insufficient depth is a primary cause of artifacts in single-step models. We therefore experiment with end-to-end training of extra-deep single-step models. In theory, extra-deep single-step models can outperform their multi-step counterparts because they can pass hidden states end-to-end without projecting into and reinterpreting from the data space, require no manual definition of timestep discretizations, and are trained without teacher forcing.
As illustrated in Figure 4, our extra-deep models use transformer block repetition (Dehghani et al., 2018). The hidden state from the initial pass is recycled. A timestep-like embedding is still provided to the transformer blocks only to distinguish repetition iterations, but the entire network is trained end-to-end for single-forward generation without any intermediate supervision. This design allows us to match the number of parameters and the model behavior of the multi-step counterpart exactly for comparison.
@@ 4.1 Setups and Training Details
We train on class-conditional ImageNet (Russakovsky et al., 2015) generation to compare with prior work. We follow standard protocols by resizing the images to $\times256$ and applying horizontal flips. We use a pre-trained variational autoencoder (VAE) [^1] https://huggingface.co/stabilityai/sd-vae-ft-mse (Rombach et al., 2022) and train the models in $32\times 32\times 4$ latent space. Evaluations use Fréchet Inception Distance on 50k class-balanced samples (FID-50k) (Heusel et al., 2017) against the entire train set.
@@
We use an initial learning rate of $1\times 10^{-4}$ and a batch size of 256 consistent with prior works (Peebles and Xie, 2023; Frans et al., 2024; Geng et al., 2025). We use the AdamW (Loshchilov and Hutter, 2017) optimizer with $\beta_{1}=0,\beta_{2}=0.9$. We set weight decay to 0.01 and EMA decay to 0.9999. We follow MeanFlow’s definition of model sizes: B/2, M/2, L/2, XL/2, where 2 denotes the patch size. $G$ and $D$ always use models of the same size. We use separate dataloaders for $G$ and $D$. Epochs are measured by the number of images seen by $G$. Since different models reach their peak FID at different epochs, we report the earliest epoch at which the best FID is reached.
Additional training techniques and details.
We provide details of additional training techniques in Appendix C, including the use of EMA reload, discriminator reset, and discriminator augmentation (DA) (Karras et al., 2020a). We provide additional training details in Appendix D, including the learning rate and OT decay.
@@
We train the model without guidance until it reaches its best FID and then continue training with guidance. Our classifier is trained from scratch on ImageNet using the cross-entropy objective for 30 epochs. It uses the same B/2 transformer architecture in latent space. We do not amortize the scale of CG into models for a fair comparison with prior works.
@@
We increase depth only in $G$ while keeping $D$ at its standard depth. The learning rate of $G$ is reduced by the repetition factor. Extra-deep models are trained end-to-end from scratch following the single-step objective.